EXERCISE 16. Urn A contains 4 white and 3 red marbles, urn B contains 2 white and 5 red marbles. A marble is to be selected from urn A and another marble is to be selected from urn B. What is the probability that the two marbles selected are white? Answer. 8/49.

Please show your workings. I have shown the possible answer already. Thank you!
Answer #1

Ans:

Urn A: 4 white and 3 red marbles

Urn B: 2 white and 5 red marbles

Probability that a white marble is selected from Urn A=4/(4+3)=4/7

Probability that a white marble is selected from Urn B=2/(2+5)=2/7

As,both events are independent,so

Probability that the two marbles selected are white=(4/7)*(2/7)=8/49